OpenStudy (anonymous):

can someone help me 3 [^3sqrt](2a)-6[^3sqrt](2a)

OpenStudy (anonymous):

\[3 \sqrt[3]{2a}-6\sqrt[3]{2a}\]

OpenStudy (kc_kennylau):

like terms :D

OpenStudy (anonymous):

can you help me understand it

OpenStudy (kc_kennylau):

what is the question anyway? simplify it?

OpenStudy (anonymous):

yes what is the simplest form

OpenStudy (kc_kennylau):

If it confuses you, let \(\sqrt[3]{2a}\) be \(b\)

OpenStudy (kc_kennylau):

\(3b-6b=?\)

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

wait no -3 am i right?

OpenStudy (kc_kennylau):

actually it's -3b lol

OpenStudy (anonymous):

ok so the answer would be \[-3\sqrt[3]{2a}\]

OpenStudy (kc_kennylau):

yep

OpenStudy (anonymous):

anything else or is that its simplest form

OpenStudy (kc_kennylau):

yep simplest form

OpenStudy (anonymous):

ok thx heres a medel

OpenStudy (kc_kennylau):

lol thanks so much :D